OpenStudy (anonymous):

If the graph of f(x) = x2 is shifted 9 units to the left, then what would be the equation of the new graph?

OpenStudy (accessdenied):

If we moved the y=x^2 at x=0 over 9 units, we'd use the value (x+9)^2 because (-9,0) is the point where the graph should touch the x-axis, which is the 0 of the new function. Not really formal, but just giving an idea. It would becmoe: y = (x+9)^2